第 8 回福岡数論研究集会

本研究集会は日本学術振興会科研費基盤 B (No.23340010) （代表者：金子昌信）の援助により開催されます。

日時：2013 年 8 月 8 日 (木) − 8 月 10 日 (土)

場所：九州大学 (伊都キャンパス) 数理学研究教育棟大講義室２福岡市西区元岡 744

プログラム

8 月 8 日 (木)

13 : 30 − 14 : 30 Jerome T. Dimabayao （九州大学）

On the vanishing of cohomologies of p-adic Galois representations associated with abelian varieties

14 : 45 − 15 : 45 小関祥康（京都大学）

On congruences of Galois representations of number fields

16 : 00 − 17 : 00 飯島優（京都大学）

A pro-l version of the congruence subgroup problem of genus one

8 月 9 日 (金)

10 : 00 − 11 : 00 三柴善範（九州大学）

Carlitz ポリログの値の代数的独立性について

11 : 15 − 12 : 15 田坂浩二（九州大学）

On a relation among totally odd triple zeta values

13 : 45 − 14 : 45 山名俊介（九州大学）

Periods of automorphic forms: the case of (GL(n, E), GL(n, F))

15 : 00 − 16 : 00 水野義紀（徳島大学）

Kernel of twisted symmetric square of elliptic modular forms

16 : 15 − 17 : 15 小松尚夫（弘前大学）

多重ベルヌーイ多項式を含むベルヌーイ多項式の積和公式

19 : 00 −  懇親会 (天神周辺を予定)

8 月 10 日 (土)

10 : 00 − 11 : 00 藤井俊（金沢工業大学）

虚アーベル体の一般 Greenberg 予想について

11 : 15 − 12 : 15 諏訪紀幸（中央大学）

Kummer theory for algebraic tori and normal basis problem: some examples

※ 研究集会の最新情報は下記アドレスでご確認ください。

  http://auemath.aichi-edu.ac.jp/~ykishi/FSNT/13-fukuokaNT.html

世話人：金子昌信（九州大学) mkaneko[at]math.kyushu-u.ac.jp
権寧魯（九州大学) ygon[at]math.kyushu-u.ac.jp
岸康弘（愛知教育大学)ykishi[at]auecc.aichi-edu.ac.jp